# Kähler compactification of $\mathbb{C}^{n}$ and Reeb dynamics

### v1

@@ 3.1. Quasi-regular Sasaki manifolds
Return to the quasi-regular Sasaki case, the contact manifold $M$ has a contact form $\eta$ whose Reeb vector field ${\mathfrak{v}}$ generates an $S^{1}$-action and $\mathcal{Y}=M/S^{1}$ is a symplectic orbifold. The standing topological assumption in this subsection is that
$c^{\mathbb{Q}}_{1}(M)=0$
which guarantees a $\mathbb{Q}$-valued indices ($\mu_{\mathrm{LCZ}},\mu_{\mathrm{RS}}$) after choosing a trivialization of $\det_{\mathbb{C}}\oplus^{N}\xi$, and is a consequence of the cone singularity being $\mathbb{Q}$-Gorenstein. In fact, it is equivalent to the singularity being $\mathbb{Q}$-Gorenstein given that the singularity is klt by [McL16, Lemma 3.3] and [BdFFU14, Corollary 5.17]. In the special case here, we do not need to assume $H^{1}(M;\mathbb{Q})=0$, as all Reeb orbits of a conical contact form are rationally null-homologous making the index independent of the trivialization. Moreover, in the question of compactification of $\mathbb{C}^{n}$, we have $H^{2}(M;\mathbb{Z})=0$ by 2.5 guaranteeing $\mathbb{Z}$-valued indices. Let $N$ be a positive integer such that $Nc_{1}(\xi)=0$. Note that for any $x\in M$, the direct sum of $N$ copies of the linearization of the $S^{1}$-action along the orbit $S^{1}\cdot x$ gives rise to a loop in $\mathrm{Sp}((2n-2)N)$ by trivializing $\oplus^{N}\xi\subset TM$ along the orbit, where the trivialization matches with the unique (by our topological assumptions) global trivialization of $\det_{\mathbb{C}}\oplus^{N}\xi$. We use $R(M)$ to denote the Maslov index $\mu$ of this loop divided by $N$, which is independent of the point $x$ by the homotopy invariant property of the Maslov index.
@@
Let $(M,\xi)$ be a contact manifold with $c_{1}^{\mathbb{Q}}(\xi)=0$ and $W$ a symplectic filling. For a rationally null-homologous Reeb orbit $\gamma$, and a bounding surface $u:S\to W$ such that $\partial u=\gamma$, using the unique trivialization of $u^{*}\det_{\mathbb{C}}TW$ over $S$, we get an integer lower Conley-Zehnder index $\mu^{u}_{\mathrm{LCZ}}(\gamma)$. Then the rational lower Conley-Zehnder index
@@
here we view $c_{1}(W)$ as in $H^{2}(W,M;\mathbb{Q})$ as $c_{1}^{\mathbb{Q}}(\xi)=0$ and $u_{*}[S]\in H_{2}(W,M;\mathbb{Z})$. Alternatively, let $A\in H_{2}(W;\mathbb{Q})$ such that $[A]=u_{*}[S]$ in $H_{2}(W,M;\mathbb{Q})$, such $A$ exists as $u_{*}[\partial S]=0\in H_{1}(M;\mathbb{Q})$. Then the rational lower Conley-Zehnder index $\mu^{\mathbb{Q}}_{\mathrm{LCZ}}(\gamma)=\mu_{\mathrm{LCZ}}^{u}(\gamma)-2\langle c_{1}(W),A\rangle$.
@@
We assume $Nc_{1}(\xi)=0$, and write $\phi(t)$ as the linearized flow, applying Property (4) to $\oplus^{N}\phi(t)$ for trivializations from $u$ and the global trivialization of $\det_{\mathbb{C}}\oplus^{N}\xi$, which are differed by $2N\langle c_{1}(W),u_{*}[S]\rangle$. ∎
Under the above setup, by the $\mathbb{Q}$-Gorenstein property or $c_{1}^{\mathbb{Q}}(\xi)=0$, there is $r(\mathcal{Y},\mathcal{L})\in\mathbb{Q}$ such that $-K_{\mathcal{Y}}=r(\mathcal{Y},\mathcal{L})\mathcal{L}$. In the case of 2.5, $r(\mathcal{Y},\mathcal{L})$ is the Fano index of $\mathcal{Y}$.
@@
We can compute the Maslov index by the induced $S^{1}$-symplectic action on the symplectization $\mathbb{R}_{s}\times M$, where the symplectic tangent bundle along an orbit is the contact structure plus $\langle{\mathfrak{v}},\partial_{s}\rangle$. Along with the obvious trivialization of $\langle{\mathfrak{v}},\partial_{s}\rangle$, the linearized action along the orbit has an extra block of $2\times 2$ identity matrix compared to the contact case. In particular, the Maslov index does not change. If we compute the Maslov index using the induced trivialization from the bounding disk $D$ in the orbifold line bundle $\mathcal{L}^{-1}$, then the loop of the symplectic matrices is just a full rotation in the two coordinates in the disk $D$, hence the Maslov index is $1$. As $12R(M)-n+$ 1 is $\mu^{\mathbb{Q}}_{\mathrm{LCZ}}$ of the principle Reeb orbit (the $S^{1}$ orbit that is not multiply covered) on $M$, $R(M)$ enjoy the same property as $\mu^{\mathbb{Q}}_{\mathrm{LCZ}}$ in 3.1 without the factor $2$. Let $A$ be a closed rational chain whose intersection number with $\mathcal{Y}$ is $1$, then $[A]=[D]$ in $H_{2}(\mathcal{L}^{-1},M;\mathbb{Q})$ and $\langle c_{1}(\mathcal{L}^{-1}),A\rangle=1$. We have $R(M)=1-\langle c_{1}(\mathcal{Y})+c_{1}(\mathcal{L}^{-1}),A\rangle=r(\mathcal{Y},\mathcal{L})$. ∎
We consider the finite set $S$ of isotropy groups (including the trivial group, which is the isotropy group for a generic point) of the $S^{1}$ action on $M$. $S$ is equipped with a partial order, we say $G_{x}>G_{y}$ if $G_{y}\subset G_{x}\subset S^{1}$ is a subgroup. For $G\in S$, the quotient of the fixed point set $M^{G}/S^{1}$ gives rise to a branch locus of the quotient Kähler orbifold $\mathcal{Y}=M/\langle{\mathfrak{v}}\rangle$ giving $M$ a stratification over the partial order set $S$. For non-minimal $G\in S$, we use $G^{-}$ to denote the unique maximal element that is smaller than $G$. We formally define $G^{-}=\emptyset$ when $G$ is the minimal element of $S$. We may rescale the contact form, such that the principle orbit (the simple Reeb orbit over a non-singular point of $M/S^{1}$) has a period $1$.
@@
    The Morse-Bott families of Reeb orbits are parameterized by $G\in S$, $k\in G\backslash G^{-}$ and $\ell\in\mathbb{Z}_{\geq 0}$, where the family of parameterized Reeb orbit is diffeomorphic to $M^{G}$ via the starting point of the Reeb orbit and the period of the orbit is $\ell+\frac{k}{|G|}$ (we parameterize $S^{1}$ by $[0,1]$). The corresponding Morse-Bott family of unparameterized Reeb orbits is diffeomorphic to the orbifold $M^{G}/S^{1}$, which is a branch locus $\mathcal{Y}_{G}$ of the base orbifold $\mathcal{Y}$ with isotropy $G$. We will call $(G,k,\ell)$ is the signature of the Reeb orbit.
@@
    The lower Conley Zehnder index $\mu_{\mathrm{LCZ}}(G,k,\ell)$ and Robbin-Salamon index $\mu_{\mathrm{RS}}(G,k,\ell)$ of orbits with signature $(G,l,\ell)$ have the following properties:
@@
      $\mu_{\mathrm{LCZ}}(G,k,\ell)=\mu_{\mathrm{RS}}(G,k,\ell)-\dim_{\mathbb{C}}{\mathcal{Y}_{G}}$;
@@
      $\mu_{\mathrm{RS}}(\{e\},0,\ell)=2\ell R(M)$;
@@
    If $R(M)\geq 0$, then
    $\inf_{\gamma}(\mathrm{lSFT}(\gamma))=\min_{G,k}\{\mu_{\mathrm{RS}}(G,k,0)-\dim_{\mathbb{C}}\mathcal{Y}_{G}+n-3\}\cup\{2R(M)-2\},$
    where $\gamma$ is a Reeb orbit of $\eta$.
@@
The description of the Morse-Bott family follows from that the Reeb flow is the $S^{1}$-action. $\mu_{\mathrm{LCZ}}(G,k,\ell)=\mu_{\mathrm{RS}}(G,k,\ell)-\dim_{\mathbb{C}}{\mathcal{Y}_{G}}$ follows from (16). Since the Robin-Salamon index for a loop of symplectic matrices is twice the Maslov index, we have $\mu_{\mathrm{RS}}(\{e\},e,\ell)=2\ell R(M)$. The linearized flow for $\mu_{\mathrm{RS}}(G,k,\ell)$ is the linearized flow for $\mu_{\mathrm{RS}}(G,k,0)$ concatenated with $\ell$-multiple of the loop of symplectic matrices from the $S^{1}$ action, then $\mu_{\mathrm{RS}}(G,k,\ell)=\mu_{\mathrm{RS}}(G,k,0)+2\ell R(M)$ by the concatenation property. To see 2(d), the mod $2$ lower Conley Zehnder index has two components: one from $M^{G}$ which is $\dim_{\mathbb{C}}\mathcal{Y}_{G}$ as it is a loop of symplectic matrices in $\mathrm{Sp}(\dim_{\mathbb{R}}\mathcal{Y}_{G})$ and one from the normal direction, which is $\frac{1}{2}(\dim M-\dim M^{G})=n-1-\dim_{\mathbb{C}}\mathcal{Y}_{G}\mod 2$ by property (5) of the Conley-Zehnder indices, as the endpoint/linearized return map is complex-linear. Then 2(d) follows from property (3). Since
$\mathrm{lSFT}(\gamma)=\mu_{\mathrm{RS}}(G,k,\ell)-\dim_{\mathbb{C}}\mathcal{Y}_{G}+n-3,$
When $R(M)\geq 0$, 2(b) implies that $\inf_{\gamma}\{\mathrm{lSFT}(\gamma)\}$ is attained by Reeb orbits with $\ell=0$, i.e.
$\inf_{\gamma}(\mathrm{lSFT}(\gamma))=\min_{G,k}\{\mu_{\mathrm{RS}}(G,k,0)-\dim_{\mathbb{C}}\mathcal{Y}_{G}+n-3\}\cup\{2R(M)-2\}.$
In the following, we compute directly that $\inf_{\gamma}\mathrm{lSFT}(\gamma)$ is twice the minimal discrepancy in (14) for the conical contact form $\eta$. We will give two proofs of (17). The one given below uses axiomatic properties of the Conley-Zehnder index. Another proof is contained in Appendix A.
@@
We consider the orbifold line bundle $\mathcal{L}^{-1}$ as a symplectic filling of the link $M$. The local symplectic model of $\mathcal{L}^{-1}$ is ×$\mathbb{C}\times\mathbb{C}^{n-1}$ with cyclic action $\frac{1}{m}(w_{1},\ldots,w_{n})$, where the first factor is fiber direction and $(m,w_{1})=1$ as the link is smooth. Let $\gamma$ denote the Reeb orbit $(e^{2\pi w_{1}\sqrt{-1}t/m},0)$ for $t\in[0,1]$. We consider the trivialization $\gamma^{*}T(\mathbb{C}\times\mathbb{C}^{n-1}/\frac{1}{m}(w_{1},\ldots,w_{n}))$,
$S^{1}\times\mathbb{C}^{n}\to\gamma^{*}T(\mathbb{C}\times\mathbb{C}^{n-1}/\frac{1}{m}(w_{1},\ldots,w_{n}))$
$(t,v)\mapsto(e^{2\pi w_{1}\sqrt{-1}t/m},0)\oplus(e^{2\pi w_{1}\sqrt{-1}t/m}v_{1},\ldots,e^{2\pi w_{n}\sqrt{-1}t/m}v_{n}).$
@@
(18) $t\to\mathrm{diag}(1,e^{2\pi(m-w_{2})\sqrt{-1}t/m},\ldots,e^{2\pi(m-w_{n})\sqrt{-1}t/m}),\quad t\in[0,1]$
Note that although we trivialize $\gamma^{*}T(\mathbb{C}\times\mathbb{C}^{n-1}/\frac{1}{m}(w_{1},\ldots,w_{n}))$ not just $\gamma^{*}\xi$, the linearized flow in the fiber direction is identity, we may compute the SFT degree from the remaining block of the linearized flow. With such trivialization $\tau$, we have
$\mathrm{lSFT}^{\tau}(\gamma)=\#\{i|i\geq 2,w_{i}>0\}+2\#\{i|i\geq 2,w_{i}=0\}-\#\{i|i\geq 2,w_{i}=0\}+n-3=2n-4,$
where #$\#\{i|i\geq 2,w_{i}>0\}$ is from those axes with $w_{i}>0$ by Property (6) of the Conley-Zehnder index, $2\#\{i|i\geq 2,w_{i}=0\}$ from those axes with $w_{i}=0$ by Property (3) and (a), and #$-\#\{i|i\geq 2,w_{i}=0\}$ is $-\dim_{\mathbb{C}}\mathcal{Y}_{\mathbb{Z}/m}=-\frac{1}{2}\dim\ker(\phi(1)-\mathrm{id})$ in (16). Similarly, we have
$\mathrm{lSFT}^{\tau^{m}}(\gamma^{m})=2\sum_{i=2}^{n}(m-w_{i})-2,$
where $\tau^{m}$ is the induced trivialization on $\gamma^{m}$, under which the linearized flow is (18) for $t\in[0,m]$. Since $\gamma^{m}$ is $w_{1}$-th cover of a principle orbit, by 3.3, $\mathrm{lSFT}$ using a global trivialization is
@@
Therefore we can compute the discrepancy between the trivialization $\tau$ and the global trivialization and conclude that:
@@
Therefore $\inf_{\gamma}\mathrm{lSFT}(\gamma)$ is the minimal discrepancy in (14). ∎

### v3

@@ 3.1. Quasi-regular Sasaki manifolds
Returning to the quasi-regular Sasaki case, the contact manifold $M$ has a contact form $\eta$ whose Reeb vector field $\mathfrak{v}$} generates an $S^{1}$-action and $\mathcal{Y}=M/S^{1}$ is a symplectic orbifold. The standing topological assumption in this subsection is that
$c^{\mathbb{Q}}_{1}(\xi)=0.$
This is a consequence of the cone singularity being $\mathbb{Q}$-Gorenstein and guarantees $\mathbb{Q}$-valued indices ($\mu_{\mathrm{LCZ}},\mu_{\mathrm{RS}}$) after choosing a trivialization of $\det_{\mathbb{C}}\oplus^{N}\xi$. In fact, by [McL16, Lemma 3.3] and [BdFFU14, Corollary 5.17], it is equivalent to the singularity being $\mathbb{Q}$-Gorenstein given that the singularity is klt. In the special case here, we do not need to assume $H^{1}(M;\mathbb{Q})=0$, as all Reeb orbits of a conic contact form are rationally null-homologous, making the index independent of the trivialization. Moreover, in the question of compactification of $\mathbb{C}^{n}$, we have $H^{2}(M;\mathbb{Z})=0$ by 2.10 guaranteeing $\mathbb{Z}$-valued indices. Let $N$ be a positive integer such that $Nc_{1}(\xi)=0$. Note that for any $x\in M$, the direct sum of $N$ copies of the linearization of the $S^{1}$-action along the orbit $S^{1}\cdot x$ gives rise to a loop in $\mathrm{Sp}((2n-2)N)$ by trivializing $\oplus^{N}\xi\subset TM$ along the orbit, where the trivialization matches with the unique (by our topological assumptions) global trivialization of $\det_{\mathbb{C}}\oplus^{N}\xi$. We use $R(M)$ to denote the Maslov index $\mu$ of this loop divided by $N$, which is independent of the point $x$ by the homotopy invariant property of the Maslov index.
@@
Let $(M,\xi)$ be a contact manifold with $c_{1}^{\mathbb{Q}}(\xi)=0$ and $W$ a symplectic filling. For a rationally null-homologous Reeb orbit $\gamma$, and a bounding surface $u:S\to W$ such that $\partial u=\gamma$, using the unique trivialization of $u^{*}\det_{\mathbb{C}}TW$ over $S$, we get an integer lower Conley-Zehnder index $\mu^{u}_{\mathrm{LCZ}}(\gamma)$. The rational lower Conley-Zehnder index is given by
@@
where we view $c_{1}(W)$ as in $H^{2}(W,M;\mathbb{Q})$ as $c_{1}^{\mathbb{Q}}(\xi)=0$ and $u_{*}[S]\in H_{2}(W,M;\mathbb{Z})$. Alternatively, if $A\in H_{2}(W;\mathbb{Q})$ satisfy $A]=u_{*}[S]$ in $H_{2}(W,M;\mathbb{Q})$ (such $A$ exists as $u_{*}[\partial S]=0\in H_{1}(M;\mathbb{Q})$), then the rational lower Conley-Zehnder index $\mu^{\mathbb{Q}}_{\mathrm{LCZ}}(\gamma)=\mu_{\mathrm{LCZ}}^{u}(\gamma)-2\langle c_{1}(W),A\rangle$.
@@
Assume $Nc_{1}(\xi)=0$, and write $\phi(t)$ as the linearized flow of the Reeb flow. We then apply Property ((P4)) to $\oplus^{N}\phi(t)$ for trivializations from $u$ and the global trivialization of $\det_{\mathbb{C}}\oplus^{N}\xi$, which are differed by $2N\langle c_{1}(W),u_{*}[S]\rangle$. ∎
Under the above setup, by the $\mathbb{Q}$-Gorenstein property or $c_{1}^{\mathbb{Q}}(\xi)=0$, there is $r(\mathcal{Y},\mathcal{L})\in\mathbb{Q}$ such that $K_{\mathcal{Y}}=r(\mathcal{Y},\mathcal{L})\mathcal{L}$. In the case of 2.10, $r(\mathcal{Y},\mathcal{L})$ is the Fano index of $\mathcal{Y}$.
@@
We can compute the Maslov index by the induced $S^{1}$-symplectic action on the symplectization $\mathbb{R}_{s}\times M$, where the symplectic tangent bundle along an orbit is the contact structure direct sum with $\langle{\mathfrak{v}},\partial_{s}\rangle$. Along with the obvious trivialization of $\langle{\mathfrak{v}},\partial_{s}\rangle$, the linearized action along the orbit has an extra block of $2\times 2$ identity matrix compared to the contact case. In particular, the Maslov index does not change. If we compute the Maslov index using the induced trivialization from the bounding disk DD–the fiber disk–in the orbifold line bundle $\mathcal{L}^{-1}$, then the loop of the symplectic matrices is just a full rotation in the two coordinates in the disk $D$, hence the Maslov index is $1$. As 2R(M)n+12R(M)-n+1 is $\mu^{\mathbb{Q}}_{\mathrm{LCZ}}$ of the principal Reeb orbit (the $S^{1}$ orbit that is not multiply covered) on $M$, $R(M)$ enjoys the same property as $\mu^{\mathbb{Q}}_{\mathrm{LCZ}}$ in 3.1 without the factor $2$. Let $A$ be a closed rational chain whose intersection number with $\mathcal{Y}$ is $1$, we may view $A$ as contained in the zero section in $\mathcal{L}^{-1}$. Then we have $[A]=[D]$ in $H_{2}(\mathcal{L}^{-1},M;\mathbb{Q})$ and $\langle c_{1}(\mathcal{L}^{-1}),A\rangle=1$. We have $R(M)=1-\langle c_{1}(Tot(\mathcal{L}^{-1})),A\rangle=1-\langle c_{1}(\mathcal{Y})+c_{1}(\mathcal{L}^{-1}),A\rangle=r(\mathcal{Y},\mathcal{L})$, where $Tot(\mathcal{L}^{-1})$ is the total space of $\mathcal{L}^{-1}$ as a complex orbifold. ∎
We consider the finite set $S$ of isotropy groups (including the trivial group, which is the isotropy group for a generic point) of the $S^{1}$ action on $M$. The set $S$ is equipped with a partial order, we say $G_{x}>G_{y}$ if $G_{y}\subset G_{x}\subset S^{1}$ is a subgroup. For $G\in S$, the quotient of the fixed point set $M^{G}/S^{1}$ gives rise to a branch locus $\mathcal{Y}_{G}$ of the quotient Kähler orbifold $\mathcal{Y}=M/\langle{\mathfrak{v}}\rangle$ giving $M$ a stratification over the partial order set $S$. For non-minimal $G\in S$, we use $G^{-}$ to denote the unique maximal element that is smaller than $G$. We formally define $G^{-}=\emptyset$ when $G$ is the minimal element of $S$. We may rescale the contact form, such that the principal orbit (the simple Reeb orbit over a non-singular point of $M/S^{1}$) has a period $1$. For $G\in S$, we write $\mathcal{Y}_{G}=\sqcup_{i\in C_{G}}\mathcal{Y}^{i}_{G}$ as the decomposition into disconnected components. $M^{G}$ has a similar decomposition into $\sqcup_{i\in C_{G}}M^{G}_{i}$
@@
[…]
@@
    The lower Conley-Zehnder index $\mu_{\mathrm{LCZ}}(G,k,\ell,i)$ and Robbin-Salamon index $\mu_{\mathrm{RS}}(G,k,\ell,i)$ of orbits with signature $(G,k,\ell,i)$ satisfy the following properties:
@@
      $\mu_{\mathrm{LCZ}}(G,k,\ell,i)=\mu_{\mathrm{RS}}(G,k,\ell,i)-\dim_{\mathbb{C}}{\mathcal{Y}^{i}_{G}}$;
@@
      $\mu_{\mathrm{RS}}(\{e\},0,\ell)=2\ell R(M)$ (there is only one component for $\pi_{0}(\mathcal{Y}_{\{e\}}=\mathcal{Y})$);
@@
    If $R(M)\geq 0$, then inf_γ(lSFT(γ)) = min_G,k,i { μ_RS(G,k,0,i)-dim_CY^i_G+n-3}∪{2R(M)-2}, where $\gamma$ ranges over Reeb orbits of $\eta$.
@@
The description of the Morse-Bott family follows from the fact that the Reeb flow is the $S^{1}$-action. The identity $\mu_{\mathrm{LCZ}}(G,k,\ell,i)=\mu_{\mathrm{RS}}(G,k,\ell,i)-\dim_{\mathbb{C}}{\mathcal{Y}^{i}_{G}}$ follows from (17). Since the Robin-Salamon index for a loop of symplectic matrices is twice the Maslov index, we have $\mu_{\mathrm{RS}}(\{e\},0,\ell)=2\ell R(M)$. The linearized flow for $\mu_{\mathrm{RS}}(G,k,\ell,i)$ is the linearized flow for $\mu_{\mathrm{RS}}(G,k,0,i)$ concatenated with $\ell$-multiple of the loop of symplectic matrices from the $S^{1}$ action, and so $\mu_{\mathrm{RS}}(G,k,\ell,i)=\mu_{\mathrm{RS}}(G,k,0,i)+2\ell R(M)$ by the concatenation property. To see 2(d), the mod $2$ lower Conley-Zehnder index is the sum of the mod 2 indices from two components, as the matrix path decomposes into two components: one from $M^{G}$ and the other from the normal direction. In the direction of $M^{G}$, the mod $2$ lower Conley-Zehnder index is $\dim_{\mathbb{C}}\mathcal{Y}^{i}_{G}$ as it is a loop of symplectic matrices in $\mathrm{Sp}(\dim_{\mathbb{R}}\mathcal{Y}^{i}_{G})$. In the normal direction, as the endpoint/linearized return map is complex-linear, we know that the parity of the complex dimension of the normal bundle, i.e. $\frac{1}{2}(\dim M-\dim M^{G}_{i})=n-1-\dim_{\mathbb{C}}\mathcal{Y}^{i}_{G}\mod 2$, is the same as the parity of the Conley-Zehnder index in the normal direction, by property ((P5)), as $\mathrm{sign}\det_{\mathbb{R}}(\mathrm{id}-\text{complex linear matrix})=1$ when it (the linearized return map in the normal direction) is non-singular. Then 2(d) follows from property ((P3)). For a Reeb orbit of signature $(G,k,\ell,i)$, we have the identity
$\mathrm{lSFT}(\gamma)=\mu_{\mathrm{RS}}(G,k,\ell,i)-\dim_{\mathbb{C}}\mathcal{Y}^{i}_{G}+n-3,$
When $R(M)\geq 0$, 2(b) implies that $\inf_{\gamma}\{\mathrm{lSFT}(\gamma)\}$ is attained by Reeb orbits with $\ell=0$ or $G=\{e\}$ and $\ell=1$, i.e.
$\inf_{\gamma}(\mathrm{lSFT}(\gamma))=\min_{G\neq\{e\},k,i}\{\mu_{\mathrm{RS}}(G,k,0,i)-\dim_{\mathbb{C}}\mathcal{Y}^{i}_{G}+n-3\}\cup\{2R(M)-2\}.$
In the following, we compute directly that $\inf_{\gamma}\mathrm{lSFT}(\gamma)$ is twice the minimal discrepancy in (15) for the conic contact form $\eta$. We will give two proofs of (19). The one given below uses axiomatic properties of the Conley-Zehnder index. Another proof is contained in Appendix A.
@@
Locally, the holomorphic vector bundle is modeled on $\mathbb{C}\times\mathbb{C}^{n-1}$ with a cyclic action $\frac{1}{m}(w_{1},\ldots,w_{n})$, where the first factor is the fiber direction and $(m,w_{1})=1$ as the link is smooth. The contact form on the associated $S^{1}$ bundle, i.e. $M$, is given by a connection form whose curvature is $\sqrt{-1}$ times the symplectic form on the base $\mathcal{Y}$. The choice of connections does not matter locally, as they differ by a gauge transformation. In the tangent space of 0 in the base $\mathbb{C}^{n-1}$, complex subspaces are symplectic, and the symplectic complement of an invariant symplectic subspace is also invariant, as the symplectic form is invariant. Hence there exists a symplectic basis such that the tangent space at 0, as a symplectic vector space, is $(\mathbb{C}^{n-1},\omega_{std})$ with a cyclic quotient action $\frac{1}{m}(w_{2},\ldots,w_{n})$. By Moser’s trick, one can argue that the symplectic form on the base is locally modeled on $(\mathbb{C}^{n-1},\omega_{std})$ with a cyclic quotient action $\frac{1}{m}(w_{2},\ldots,w_{n})$. As $\omega_{std}=\mathrm{d}\lambda_{std}=\mathrm{d}(\frac{1}{2}\sum_{i=2}^{n}(x_{i}\mathrm{d}y_{i}-y_{i}\mathrm{d}x_{i}))$ and $\lambda_{std}$ is invariant under the cyclic action, we can use $\mathrm{d}\theta+\lambda_{std}$ as the connection on the trivialization. Therefore the contact form is locally modeled on $(S^{1}\times\mathbb{C}^{n-1},\mathrm{d}\theta+\lambda_{std})$ with a cyclic action action $\frac{1}{m}(w_{1},\ldots,w_{n})$. Note that we are only claiming the contact form, not the Sasaki structure, is standard, as Moser’s trick only preserves the symplectic structure. With such a standard contact model, the Reeb flow is given by $\partial_{\theta}$. The projection $\pi:\xi=\ker(\mathrm{d}\theta+\lambda_{std})\subset T(S^{1}\times\mathbb{C}^{n-1})\to\mathbb{C}^{n-1}$ is an isomorphism, and $\pi_{*}(\mathrm{d}(\theta+\lambda_{std}))=\mathrm{d}\lambda_{std}=\omega_{std}$ on $\mathbb{C}^{n-1}$. In order to make the principal orbits have period $1$ as in 3.3, we rescale the contact form to $\frac{1}{2\pi}(\mathrm{d}\theta+\lambda_{std})$, such that the Reeb flow is $\phi_{t}(\theta,v)=(\theta+2\pi t,v)$. Let $\gamma$ denote the Reeb orbit $(e^{2\pi w_{1}\sqrt{-1}t/m},0)$ for $t\in[0,1]$, we consider the symplectic trivialization $\gamma^{*}\xi$ by
$S^{1}\times\mathbb{C}^{n-1}\to\gamma^{*}\xi$
$(t,v=(v_{2},\ldots,v_{n}))\mapsto\pi^{-1}(e^{2\pi w_{2}\sqrt{-1}t/m}v_{2},\ldots,e^{2\pi w_{n}\sqrt{-1}t/m}v_{n})$
@@
(20) $t\to\mathrm{diag}(e^{-2\pi w_{2}\sqrt{-1}t/m},\ldots,e^{-2\pi w_{n}\sqrt{-1}t/m}),\quad t\in[0,1]$
With such a trivialization $\tau$, we have
$\mathrm{lSFT}^{\tau}(\gamma)=-\#\{i|i\geq 2,w_{i}>0\}-\#\{i|i\geq 2,w_{i}=0\}+n-3=-2,$
where $\{i|i\geq 2,w_{i}>0\}$ is from those axes with $w_{i}>0$ by Property ((P6)) of the Conley-Zehnder index, and $\{i|i\geq 2,w_{i}=0\}$ is $\dim_{\mathbb{C}}\mathcal{Y}_{\mathbb{Z}/m}=-\frac{1}{2}\dim\ker(\phi(1)-\mathrm{id})$ in (17), which is the lower Conley-Zehnder index from those axes with $w_{i}=0$ by Properties ((P7)) and ((P8)). Then the claim follows from Property ((P3)). Similarly, by Properties ((P7)) and ((P8)), we have
$\mathrm{lSFT}^{\tau^{m}}(\gamma^{m})=-2\sum_{i=2}^{n}w_{i}-2,$
where $\tau^{m}$ is the induced trivialization on $\gamma^{m}$, under which the linearized flow is (20) for $t\in[0,m]$. Since $\gamma^{m}$ is $w_{1}$-th cover of a principal orbit, by 3.3, $\mathrm{lSFT}$ using a global trivialization is
@@
Therefore, we can compute the discrepancy between the trivialization $\tau$ and the global trivialization and conclude that:
@@
Therefore $\inf_{\gamma}\mathrm{lSFT}(\gamma)$ is twice the minimal discrepancy in (15). ∎